R/3 -4=2 then what does r equal.
1
Expert's answer
2014-10-13T07:39:35-0400
R/3-4=2, add 4 to both sides, which gives R/3=2+4, R/3=6, multiply both sides by 3 and obtain R=18. Answer: R=18.